What will be the change of temperature for 300 gr of iron (Fe) which have gained 2000 cal?


1
Expert's answer
2021-02-04T14:38:34-0500
"Q=mc\Delta T,""\Delta T=\dfrac{Q}{mc}=\dfrac{2000\ cal}{300\ g\cdot 0.107\ \dfrac{cal}{g \cdot^{\circ}C}}=62.3\ ^{\circ}C."